On a map the distance from Sacramento to San francisco is 8cm. THe legend gives the actual distance as 96 miles. ON the same map,On the same map it is also 7cm from Sacramento to Modesto. What is the actual distance?
Answer by nyc_function(2741)

Le m = actual distance in miles
8cm/7cm = 96 miles/m
8/7 = 96/m
96 * 7 = 8m
672 = 8m
672/8 = m
84 = m
The actual distance is 84 miles.